# bibliography.bib

### v1

@@ A.1 Conditional average treatment effect (CATE)
Estimating the CATE has received a lot of attention in the machine learning literature (e.g.,[chipman2010bart, curth2021nonparametric, curth2021inductive, johansson2016learning, johansson2018learning, kunzel2019metalearners, kennedy2023towards, ma2025diffusion, ma2024diffpo, ma2025foundation, nie2021quasi, shalit2017estimating, wager2018estimation]). A prominent approach to CATE estimation involves meta-learners – i.e., flexible strategies that decouple treatment effect estimation from the choice of base machine learning model. First systematized by [kunzel2019metalearners], these methods have since been extended through theoretical analysis [kennedy2022minimax, nie2021quasi] and improved pseudo-outcome constructions [kennedy2023towards].
@@
As the latter is based on the doubly-robust AIPW estimator, it is hence unbiased if either propensity score or outcome regressions are correctly specified [kennedy2023towards].
@@ A.2 Causal inference with text data
Causal inference with text data has emerged as an important research direction (e.g. [veitch2020adapting, keith2020text, roberts2020adjusting, mozer2020matching, zhang2020quantifying, maiya2021causalnlp, gultchin2021operationalizing, keith2021text, daoud2022conceptualizing, zeng2022uncovering, deshpande2022deep, egami2022make, chen2024proximal]), with text variables serving multiple roles in causal frameworks.
Prior research has considered text to have various causal roles, with text serving as a treatment, mediator, outcome, or confounder. For example, previous works view text as treatment including [[, e.g.,]]tan2014effect, wood2018challenges, wang2019words, pryzant2021causal, maiya2021causalnlp, egami2022make; works view text as mediator including[[, e.g.,]] zhang2020quantifying, veitch2020adapting, gultchin2021operationalizing, keith2021text); and works view text as outcome including [[, e.g.,]]gill2015judicial, sridhar2019estimating, koroleva2019measuring). There are many works that have explored text as a confounder [[, e.g.,]]keith2020text, roberts2020adjusting, mozer2020matching, daoud2022conceptualizing, zeng2022uncovering, chen2024proximal). [keith2020text] surveyed methods leveraging text to remove confounding and challenges. Some work attempts to mitigate confounding bias by viewing text features act as proxies for unobserved confounders [veitch2020adapting, roberts2020adjusting, deshpande2022deep, chen2024proximal].
@@ A.3 NLP and language models in medical decision-making
Natural language processing (NLP) has emerged as an important tool in healthcare, enabling the extraction of meaningful information from unstructured clinical texts [[, e.g.,]] demner2009can, chen2020trends, huang2019clinicalbert, zhu2020using, wu2020deep, salunkhe2021machine, hiremath2022enhancing, hossain2023natural,hossain2024natural). Early work focused on using NLP to process clinical notes and symptom descriptions for some basic analysis and classification tasks, while more recent approaches leverage deep learning for learning the representations, particularly transformer-based models like BERT and GPT [devlin2018bert, huang2019clinicalbert, brown2020language, dhawan2024end, lahat2024assessing]. Work in this area has focused on learning the representations; for instance,, [huang2019clinicalbert] improved on previous clinical text processing methods and proposed a pretrained model on clinical notes to learn better representation for prediction. Other work used LLMs to impute the missing data in clinical notes [dhawan2024end].
Large language models (LLMs) have shown promise in medical applications, such as clinical text summarization, diagnosis, and treatment recommendations [thirunavukarasu2023large, nazi2024large, spitzer2025effect, ullah2024challenges, dhawan2024end]. For example, [kimmdagents] introduce a multi-agent framework that leverages LLMs to emulate the hierarchical diagnosis procedures, targeted for symptom classification and risk stratification tasks. However, unstructured text data often lacks explicit information about key confounders, which is critical for accurate causal inference. most NLP models, including LLMs, are designed for associative learning rather than causal reasoning [jin2023cladder, jin2023can]. Their use in causal inference, particularly in addressing text confounding, remains underexplored.
@@ A.4 Counterfactual predictions under runtime confounding
Our setting is also related to [coston2020counterfactual], which refers to a scenario where historical data contains all the relevant information needed for decision-making. Only a subset of covariates $V\subseteq X$ is available to use at runtime. The unobserved set $X\setminus V$ should include all the hidden confounders at the runtime, denoted as $Z$. They show this can induce considerable bias in the resulting prediction model when the discarded features are significant confounders. However, this setting is different from ours. They require the set $V$ at runtime known and fixed. This is a strong and unrealistic assumption in our setting. Our setting allows $V$ to be an arbitrary subset of $X$. Moreover, their setting cannot handle the modality of data, which involves structured clinical tabular data at training time and only self-reported description-based text data at inference time.
@@ B.1 Pointwise confounding bias of the naïve baseline
We give the proof by following [coston2020counterfactual].
@@ B.3 Double robustness property of the estimator
We give the proof following [kennedy2023towards]. We treat the two cases separately.
@@
$-\frac{1}{1-\pi^{x}(x)}\,\mathbb{E}\Bigl{[}(1-A)\Bigl{(}Y-\mu_{0}^{x}(x)-\delta_{0}(x)\Bigr{)}\mid X=x\Bigr{]}=\delta_{0}(x).$ (36)
@@ Appendix C Implementation details
Text generation: Given structured clinical confounders $X$, we generate the induced text confounder $\tilde{T}$ using gpt-4o-mini-2024-07-18 through the OpenAI API. For each patient, we construct prompts by templating the key features of $X$ into natural language constraints. Generated texts are approximately $150-200$ tokens.
@@
Model architecture and training: Propensity scores $\hat{\pi}^{x}(x)$ are estimated via logistic regression The CATE predictor is a 3-layer MLP with ReLU activation, batch normalization, and 0.3 dropout. We use Adam optimizer with learning rate $55e-$ 5. Models are trained for $100$ epochs on IST and $150$ on MIMIC-III with a batch size of $512$. We apply label smoothing $\alpha=0.1$ and gradient clipping (max norm=$1.0$)
@@
Baselines: We follow the implementation from https://github.com/AliciaCurth/CATENets/tree/main for most of the CATE estimators, including S-Net [kunzel2019metalearners], T-Net [kunzel2019metalearners], TARNet [shalit2017estimating], CFRNet [shalit2017estimating]. Regarding the proxy-based methods [veitch2020adapting, chen2024proximal], we follow the implementation from https://github.com/rpryzant/causal-bert-pytorch and the implementation from https://github.com/jacobmchen/proximal_w_text.
@@ D.1 The International Stroke Trial database
The International Stroke Trial (IST) [sandercock2011international] was conducted between 1991 and 1996. It is one of the largest randomized controlled trials in acute stroke treatment. The dataset comprises $19,435$ patients from $467$ hospitals across $36$ countries, enrolled within $48$ hours of stroke onset. The treatment assignments include (trial aspirin allocation and trial heparin allocation with levels for medium dose, low dose, and no heparin). The outcome is whether a patient will experience stroke again. Covariates include age, sex, presence of atrial fibrillation, systolic blood pressure, infarct visibility on CT, prior heparin use, prior aspirin use, and recorded deficits of different body parts.
@@ D.2 MIMIC-III dataset
The Medical Information Mart for Intensive Care (MIMIC-III) [johnson2016mimic] is a large, single-center database comprising information relating to patients admitted to critical care units at a large tertiary care hospital. MIMIC-III contains $38,597$ distinct adult patients. We follow the standardized preprocessing pipeline [wang2020mimic] of the MIMIC-III dataset. We use demographic variables such as gender and age, along with clinical variables including vital signs and laboratory test results upon admission, as confounders. These include glucose, hematocrit, creatinine, sodium, blood urea nitrogen, hemoglobin, heart rate, mean blood pressure, platelets, respiratory rate, bicarbonate, red blood cell count, and anion gap, among others. The binary treatment is mechanical ventilation. The outcome is the number of days a patient needs to stay in the hospital.
We follow previous work [kunzel2019metalearners, shalit2017estimating, curth2021nonparametric, curth2021inductive, kennedy2023towards] to simulate treatment and outcome as follows
@@ D.4 Experiments with varying prompt strategies
Appendix E Further discussion
Our framework leverages LLM-derived text to capture an induced text confounder during training. Hence, we acknowledge risks from LLM use, such as biases or data misrepresentation. We thus advocate for a rigorous validation of LLM outputs with domain experts to ensure reliability and safety, particularly in high-stakes contexts in medicine. Further, LLMs should be used transparently and ethically [feuerriegel2025using], with safeguards against amplifying societal biases. Future work should explore integrating uncertainty quantification to further increase trust in our TCA during real-world deployment.

### v2

@@ A.1 Conditional average treatment effect (CATE)
Estimating the CATE has received a lot of attention in the machine learning literature (e.g.,[4, 6, 7, 28, 29, 39, 33, 41, 42, 43, 47, 54, 62]). A prominent approach to CATE estimation involves meta-learners – i.e., flexible strategies that decouple treatment effect estimation from the choice of base machine learning model. First systematized by [39], these methods have since been extended through theoretical analysis [34, 47] and improved pseudo-outcome constructions [33].
@@
As the latter is based on the doubly-robust AIPW estimator, it is hence unbiased if either propensity score or outcome regressions are correctly specified [33].
@@ A.2 Causal inference with text data
Causal inference with text data has emerged as an important research direction (e.g. [61, 31, 50, 45, 72, 44, 17, 32, 8, 70, 10, 13, 2]), with text variables serving multiple roles in causal frameworks.
Prior research has considered text to have various causal roles, with text serving as a treatment, mediator, outcome, or confounder. For example, previous works view text as treatment including [[, e.g.,]]tan2014effect, wood2018challenges, wang2019words, pryzant2021causal, maiya2021causalnlp, egami2022make; works view text as mediator including[[, e.g.,]] zhang2020quantifying, veitch2020adapting, gultchin2021operationalizing, keith2021text); and works view text as outcome including [[, e.g.,]]gill2015judicial, sridhar2019estimating, koroleva2019measuring). There are many works that have explored text as a confounder [[, e.g.,]]keith2020text, roberts2020adjusting, mozer2020matching, daoud2022conceptualizing, zeng2022uncovering, chen2024proximal). [31] surveyed methods leveraging text to remove confounding and challenges. Some work attempts to mitigate confounding bias by viewing text features act as proxies for unobserved confounders [61, 50, 10, 2].
@@ A.3 NLP and language models in medical decision-making
Natural language processing (NLP) has emerged as an important tool in healthcare, enabling the extraction of meaningful information from unstructured clinical texts [[, e.g.,]] demner2009can, chen2020trends, huang2019clinicalbert, zhu2020using, wu2020deep, salunkhe2021machine, hiremath2022enhancing, hossain2023natural,hossain2024natural). Early work focused on using NLP to process clinical notes and symptom descriptions for some basic analysis and classification tasks, while more recent approaches leverage deep learning for learning the representations, particularly transformer-based models like BERT and GPT [11, 25, 1, 12, 40]. Work in this area has focused on learning the representations; for instance,, [25] improved on previous clinical text processing methods and proposed a pretrained model on clinical notes to learn better representation for prediction. Other work used LLMs to impute the missing data in clinical notes [12].
Large language models (LLMs) have shown promise in medical applications, such as clinical text summarization, diagnosis, and treatment recommendations [59, 46, 56, 60, 12]. For example, [36] introduce a multi-agent framework that leverages LLMs to emulate the hierarchical diagnosis procedures, targeted for symptom classification and risk stratification tasks. However, unstructured text data often lacks explicit information about key confounders, which is critical for accurate causal inference. most NLP models, including LLMs, are designed for associative learning rather than causal reasoning [27, 26]. Their use in causal inference, particularly in addressing text confounding, remains underexplored.
@@ A.4 Counterfactual predictions under runtime confounding
Our setting is also related to [5], which refers to a scenario where historical data contains all the relevant information needed for decision-making. Only a subset of covariates $V\subseteq X$ is available to use at runtime. The unobserved set $X\setminus V$ should include all the hidden confounders at the runtime, denoted as $Z$. They show this can induce considerable bias in the resulting prediction model when the discarded features are significant confounders. However, this setting is different from ours. They require the set $V$ at runtime known and fixed. This is a strong and unrealistic assumption in our setting. Our setting allows $V$ to be an arbitrary subset of $X$. Moreover, their setting cannot handle the modality of data, which involves structured clinical tabular data at training time and only self-reported description-based text data at inference time.
@@ B.1 Pointwise confounding bias of the naïve baseline
We give the proof by following [5].
@@ B.3 Double robustness property of the estimator
We give the proof following [33]. We treat the two cases separately.
@@
$\frac{1}{1-\pi^{x}(x)}\,\mathbb{E}\Bigl[(1-A)\Bigl(Y-\mu_{0}^{x}(x)-\delta_{0}(x)\Bigr)\mid X=x\Bigr]=\delta_{0}(x).$ (36)
@@ Appendix C Implementation details
Text generation: Given structured clinical confounders $X$, we generate the induced text confounder $\tilde{T}$ using GPT-4o mini through the OpenAI API. For each patient, we construct prompts by templating the key features of $X$ into natural language constraints. Generated texts are approximately 150200150-200 tokens.
@@
Model architecture and training: Propensity scores $\hat{\pi}^{x}(x)$ are estimated via logistic regression The CATE predictor is a 3-layer MLP with ReLU activation, batch normalization, and 0.3 dropout. We use Adam optimizer with learning rate 5e55e-5. Models are trained for $100$ epochs on IST and $150$ on MIMIC-III with a batch size of $512$. We apply label smoothing $\alpha=0.1$ and gradient clipping (max norm=$1.0$)
@@
Baselines: We follow the implementation from https://github.com/AliciaCurth/CATENets/tree/main for most of the CATE estimators, including S-Net [39], T-Net [39], TARNet [54], CFRNet [54]. Regarding the proxy-based methods [61, 2], we follow the implementation from https://github.com/rpryzant/causal-bert-pytorch and the implementation from https://github.com/jacobmchen/proximal_w_text.
@@ D.1 The International Stroke Trial database
The International Stroke Trial (IST) [53] was conducted between 1991 and 1996. It is one of the largest randomized controlled trials in acute stroke treatment. The dataset comprises $19,435$ patients from $467$ hospitals across $36$ countries, enrolled within $48$ hours of stroke onset. The treatment assignments include (trial aspirin allocation and trial heparin allocation with levels for medium dose, low dose, and no heparin). The outcome is whether a patient will experience stroke again. Covariates include age, sex, presence of atrial fibrillation, systolic blood pressure, infarct visibility on CT, prior heparin use, prior aspirin use, and recorded deficits of different body parts.
@@ D.2 MIMIC-III dataset
The Medical Information Mart for Intensive Care (MIMIC-III) [30] is a large, single-center database comprising information relating to patients admitted to critical care units at a large tertiary care hospital. MIMIC-III contains $38,597$ distinct adult patients. We follow the standardized preprocessing pipeline [63] of the MIMIC-III dataset. We use demographic variables such as gender and age, along with clinical variables including vital signs and laboratory test results upon admission, as confounders. These include glucose, hematocrit, creatinine, sodium, blood urea nitrogen, hemoglobin, heart rate, mean blood pressure, platelets, respiratory rate, bicarbonate, red blood cell count, and anion gap, among others. The binary treatment is mechanical ventilation. The outcome is the number of days a patient needs to stay in the hospital.
We follow previous work [39, 54, 6, 7, 33] to simulate treatment and outcome as follows
@@ D.4 Experiments with varying prompt strategies
Appendix E Additional acknowledge
We have access to ChatGPT Edu through our university, which is SOC 2 Type II compliant (https://openai.com/index/introducing-chatgpt-edu/). This gives us enhanced privacy controls not available in the standard API, free version, or even ChatGPT Team.
As part of this setup, we have applied for a Business Associate Agreement (BAA) (https://help.openai.com/en/articles/8660679-how-can-i-get-a-business-associate-agreement-baa-with-openai-for-the-api-services) to gain access to Zero Data Retention (ZDR). With ZDR enabled, the store parameter for /v1/responses and v1/chat/completions will always be treated as false https://platform.openai.com/docs/guides/your-data#zero-data-retention. In this setup, OpenAI does not retain any request or response data after processing, meaning there is no storage, no logging, and, more importantly, no human review. This configuration provides a privacy approach that is functionally equivalent to the one recommended by PhysioNet https://physionet.org/news/post/gpt-responsible-use, where Azure OpenAI is used with human review explicitly disabled. We confirm that our analysis is thus HIPAA-compliant and thus in line with the PhysioNet rules.
Appendix F Further discussion
Our framework leverages LLM-derived text to capture an induced text confounder during training. Hence, we acknowledge risks from LLM use, such as biases or data misrepresentation. We thus advocate for a rigorous validation of LLM outputs with domain experts to ensure reliability and safety, particularly in high-stakes contexts in medicine. Further, LLMs should be used transparently and ethically [15], with safeguards against amplifying societal biases. Future work should explore integrating uncertainty quantification to further increase trust in our TCA during real-world deployment.